OpenStudy (anonymous):

What is the slope of the line defined by the equation 2x + 3y = 18?

OpenStudy (texaschic101):

you need to put the equation in y = mx + b form and m is your slope. Can you do that ?

OpenStudy (anonymous):

is it 2

OpenStudy (raden):

alternative : the slope of line ax + by = c is m = -a/b with m = slope

OpenStudy (anonymous):

i still dont get it

OpenStudy (raden):

2x + 3y = 18, it means a = 2 and b = 3. so m = -a/b = ....

OpenStudy (anonymous):

2/18

OpenStudy (texaschic101):

try it this way....put the equation in y = mx + b form. 2x + 3y = 18 (subtract 2x from both sides) 3y = -2x + 18 (now divide by 3) y = -2/3x + 6 slope in y = mx + b form is in the m position. So what is your slope ?

OpenStudy (anonymous):

me and someone else just got 2. i dont get how to do this math.

OpenStudy (texaschic101):

did you put it in y = mx + b form ?

OpenStudy (texaschic101):

your probably just looking at the number in front of x without rearranging the equation.